\begin{thebibliography}{width}
	\bibitem{AB}
	Akta\c{s} \.{I}. and Baricz \'A., Bounds for radii of starlikeness of some $q-$Bessel functions, {\it Results Math}, \textbf{72}(1): 947--963, 2017.
	
	\bibitem{ABS}
	Akta\c{s} \.{I}., Baricz \'A. and Singh S., Geometric and monotonic properties of hyper-Bessel functions, {\it Ramanujan J.}, 1--21, doi: 10.1007/s11139-018-0105-9, 2019.
	
	\bibitem{ATO}
	Akta\c{s} \.{I}., Toklu E. and Orhan H., Radii of uniform convexity of some special functions. {\it Turk J Math}, \textbf{42}(6): 3010--3024, 2018.
	
	\bibitem{AJR}
	Ali R. M., Jain N. K. and Ravichandran V., Radii of starlikeness associated with the lemniscate of Bernoulli and the left-half plane, {\it Appl. Math. Comput.}, \textbf{218} (2012), no. 11, 6557–6565.
	
	\bibitem{AP}
	Baricz \'A. and Prajapati A. Radii of starlikeness and convexity of generalized Mittag-Leffler functions. arXiv preprint \href{https://arxiv.org/abs/1901.04333}{arXiv:1901.04333}, 2019.
	
	\bibitem{BS}
	Baricz \'{A}. and Sz\'{a}sz R. The radius of convexity of normalized Bessel functions of the first kind, {\em Anal. Appl.}, 12(5), 485--509, 2014.
	
	\bibitem{BTK}
	Baricz \'A., Toklu E. and Kad{\i}o\u{g}lu E., Radii of starlikeness and convexity of Wright functions. {\it Math. Commun.}, 23: 97--117, 2018.
	
	\bibitem{CR}
	Chaggara  H. and Romdhane N.B., On the zeros of the hyper-Bessel function, {\it Integr. Transf. Spec. Funct.}, \textbf{26}(2), 96--101, 2015.
	
	\bibitem{Goodman}
	Goodman A. W. , {\it Univalent functions. Vol. I}, Mariner Publishing Co., Inc., Tampa, FL, 1983.
	
	\bibitem{Janowski}
	Janowski W., Extremal problems for a family of functions with positive real part and for some related families, {\it Ann. Polon. Math.}, \textbf{23} (1970/1971), 159–177.
	
	\bibitem{MKR1}
	Madaan V., Kumar A., and Ravichandran V., Lemniscate Convexity and Other Properties of
	Generalized Bessel Functions. arXiv preprint \href{https://arxiv.org/abs/1902.04277}{arXiv:1902.04277}, 2019.
	
	\bibitem{MKR2}
	Madaan V., Kumar A., and Ravichandran V., Radii of starlikeness and convexity of Bessel functions. arXiv preprint \href{https://arxiv.org/abs/1906.05547}{arXiv:1906.05547}, 2019.
	
	\bibitem{mainardi}
	R. Gorenflo, Y. Luchko and F. Mainardi, Analytical properties and applications of the Wright function, {\it Fract. Calc. Appl. Anal.} \textbf{2}(4), 383--414, 1999.
	
	\bibitem{SS}
	Sok\'ol J. and Stankiewicz J., Radius of convexity of some subclasses of strongly starlike functions, {\it Zeszyty Nauk. Politech. Rzeszowskiej Mat.}, No. \textbf{19}, 101–105, 1996.
	
	\bibitem{Toklu}
	Toklu E., Radii of starlikeness and convexity of  $q-$Mittag-Leffler functions. {\it Turk J Math}, \textbf{43}(5): 2610--2630, 2019.
	
	\bibitem{TAO}
	Toklu E., Akta\c{s} \.{I}. and Orhan H., Radii problems for normalized $q-$Bessel and Wright functions. {\it Acta Univ Sapientiae Mathematica}, \textbf{11}(1): 193--213, 2019.
	
	\bibitem{VR}
	Verma S. and Ravichandran V., Radius problems for ratios of Janowski starlike functions with their derivatives, {\it Bull. Malays. Math. Sci. Soc.}, \textbf{40} (2017), no. 2, 819–840, 2017.
	
	\bibitem{wright}
	\textsc{E.M. Wright}, On the coefficients of power series having exponential singularities, {\em J. Lond. Math. Soc}, 71--79, 1933.
\end{thebibliography}

\documentclass[11 pt]{amsart}
\usepackage{amsfonts}
\usepackage{ifthen}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{graphicx}
\usepackage{amscd,amssymb,amsthm}
\usepackage{graphicx}
\usepackage{epstopdf}
\usepackage{hyperref}
\usepackage{cleveref}
\usepackage{cite}
\usepackage{mathrsfs}

\newcounter{minutes}
\divide\time by 60
\newcounter{hours}
\multiply\time by 60 \addtocounter{minutes}{-\time}

\keywords{Wright function; Lemniscate of Bernoulli; Janowski function.}
\subjclass[2010]{30C45, 30C15, 33C10}

\begin{document}
	
	\title{On some geometric properties of normalized Wright functions}

	\author[E. Toklu]{Evr{\.I}m Toklu}
	\address{Department of Mathematics, Faculty of Education, A\u{g}r{\i} {\.I}brahim \c{C}e\c{c}en University, A\u{g}r{\i}, Turkey} 
	\email{evrimtoklu@gmail.com}
	
	\author[N. Karag\"{o}z]{Nesl{\.I}han Karag\"{O}z}
	\address{Institute of Science and Technology, A\u{g}r{\i} {\.I}brahim \c{C}e\c{c}en University, A\u{g}r{\i}, Turkey} 
	\email{nslhnkrgz@gmail.com}
	
	\def\thefootnote{}
	\footnotetext{ \texttt{File:~\jobname .tex,
			printed: \number\year-0\number\month-\number\day,
			\thehours.\ifnum\theminutes<10{0}\fi\theminutes}
	} \makeatletter\def\thefootnote{\@arabic\c@footnote}\makeatother
	
	\maketitle
	
\begin{abstract}
The main purpose of the present paper is to determine the radii of starlikeness and convexity associated with lemniscate of Bernoulli and the Janowski function, $(1+Az)/(1+Bz)$ for $-1\leq B<A\leq 1,$ of normalized Wright functions. The key tools in the proof of our main results are the infinite product representation of Wright function and properties of real zeros of the Wright function and its derivative.
\end{abstract}
\section{\bf Introduction and the main results}
Let $\mathbb{D}_{r}$ be the open disk $\left\lbrace z\in\mathbb{C}:\left| z\right| <r\right\rbrace $ with the radius $r>0$ and let $\mathbb{D}=\mathbb{D}_{1}.$ Let $f:\mathbb{D}_r\to\mathbb{C}$ be the function defined by
\begin{equation}
f(z)=z+\sum_{n\geq 2}a_{n}z^{n},  \label{W0}
\end{equation}
where $r$ is less or equal than the radius of convergence of the above power series. Let $\mathcal{A}$ be the class of analytic functions of the form \eqref{W0}, that is, normalized by the conditions $f(0)=f^{\prime}(0)-1=0.$ Let $\mathcal{S}$ denote the class of functions belonging to $\mathcal{A}$ which are of univalent in $\mathbb{D}_{r}$. A function $f\in \mathcal{A}$ is said to be starlike function if $f(\mathbb{D})$ is starlike domain with the respect to the origin. It is well known fact that various subclasses of starlike function can be unified by making use of the concept of subordination. A function $f\in \mathcal{A}$ is subordinate to a function $g\in \mathcal{A},$ written as $f(z)\prec g(z),$ if there exist a Schwarz function $w$ with $w(0)=0$ and $\left| w(z)\right| <1$ such that $f(z)=g(w(z)).$ In addition, we know that if $g$ is a univalent function, then $f(z)\prec g(z)$ if and only if $f(0)=g(0)$ and $f(\mathbb{D})\subset g(\mathbb{D}).$ For an analytic function $\varphi,$ let $\mathcal{S}^{\star}(\varphi)$ denote the class of all analytic functions satisfying $1+zf'(z)/f(z)\prec \varphi(z).$ By $\mathcal{K}(\varphi)$ we mean the class of all analytic functions satisfying $1+zf''(z)/f'(z)\prec \varphi(z).$ It is worth mentioning that these classes include respectively several famous subclasses of starlike and convex functions. For instance, the class $\mathcal{S}^{\star}_{\mathcal{L}}:=\mathcal{S}^{\star}(\sqrt{1+z})$ denotes the class of lemniscate starlike functions introduced and investigated by Sok\'ol and Stankiewich \cite{SS} and the class $\mathcal{K}_{\mathcal{L}}:=\mathcal{K}(\sqrt{1+z})$ represents the class  of lemniscate convex functions. Moreover, for $-1\leq B<A\leq 1,$ the class $\mathcal{S}^{\star}[A,B]:=\mathcal{S}^{\star}((1+Az)/(1+Bz))$ is the class of Janowski starlike functions and $\mathcal{K}[A,B]:=\mathcal{K}((1+Az)/(1+Bz))$ is the class of Janowski convex functions \cite{Janowski}.

Given a class of functions $\mathcal{M}\subset\mathcal{A}$ and a function $f\in \mathcal{A},$ the $\mathcal{M}-$radius of the function $f$ is the largest number $r$ with $0\leq r \leq 1$ such that $f_{r}\in \mathcal{M},$ where $f_{r}(z):=f(rz)/r.$ If we choose $\mathcal{M}=\mathcal{S}^{\star}_{\mathcal{L}},$ the $\mathcal{M}-$radius of the function $f$, which is represented by $r^{\star}_{\mathcal{L}}(f),$ is called the radius of lemniscate starlikeness. It is indeed the largest $r$ with $0\leq r \leq 1$ such that
\[\left|\left(\frac{zf'(z)}{f(z)} \right)^2-1  \right|<1 \quad (\left| z\right|<r ).\]
If we choose $\mathcal{M}=\mathcal{K}_{\mathcal{L}},$ the $\mathcal{M}-$radius of the function $f$, which is represented by $r^{c}_{\mathcal{L}}(f),$ is called as the radius of lemniscate convexity. It is indeed the largest $r$ with $0\leq r \leq 1$ such that
\[\left|\left(1+\frac{zf''(z)}{f'(z)} \right)^2-1  \right|<1 \quad (\left| z\right|<r ).\]
If we take $\mathcal{M}=\mathcal{S}^{\star}[A,B]$ or $\mathcal{M}=\mathcal{K}[A,B],$ the respective $\mathcal{M}-$radii, which are represented by $r^{\star}_{A,B}(f)$ and $r^{c}_{A,B}(f),$ are called as the radii of Janowski starlikeness and Janowski convexity. These are respectively the largest  $r$ with $0\leq r \leq 1$ such that
\[ \left|\frac{\left( zf'(z)/f(z)\right)-1 }{A-Bzf'(z)/f(z)} \right|<1 \text{ \ and \ }      \left|\frac{zf''(z)/f'(z) }{A-B\left( 1+zf''(z)/f'(z)\right) }\right| <1 \quad (\left| z\right|<r ).\]

Recently, there has been a vivid interest on some geometric properties such as univalency,
starlikeness, convexity and uniform convexity of various special functions such as hyper-Bessel, Wright, $q-$Bessel and Mittag-Leffler functions (see \cite{ABskecul,ABS,ATO,AP,BTK,TAO}). Fore more details on the radius problems, one may consult on the studies \cite{AJR,Goodman,MKR1,Toklu,VR}. Moreover, in \cite{MKR2} the authors determined the radii of starlikeness and convexity associated with lemniscate of Bernoulli and the Janowski function $(1+Az)/(1+Bz)$.  Motivated by the above series of papers on geometric properties of special functions, in this paper our aim is to determine the radii of lemniscate starlikeness, lemniscate convexity, Janowski starlikeness and Janowski convexity of normalized Wright functions.

\end{document}